Shadee Corp. expects to sell 600 sun visors in May and 360 in June. Each visor sells for $18. Shadee’s beginning and ending finished goods inventories for May are 75 and 55 units, respectively. Ending finished goods inventory for June will be 65 units.

It expects the following unit sales for the third quarter:      

July 560
August 480
September 460


Sixty percent of Shadee’s sales are cash. Of the credit sales, 52 percent is collected in the month of the sale, 37 percent is collected during the following month, and 11 percent is never collected.

Required:

Calculate Shadee’s total cash receipts for August and September. (Do not round your intermediate calculations. Round your answers to the nearest whole dollar.)

August?

September?
Answer #1
August September
Cash sale 480*18*60% = $5184 460*18*60% = $4968
Cash received from August sale 480*18*40%*52% = 1797.12 480*18*40%*37% = 1278.72
Cash received from September sale 460*18*40%*52% = 1722.24
Cash received from July sale 560*18*40%*37% = 1491.84
Total cash receipts $8473 $7969